1. What mass of Al will react with 150g of Fe2O3? 
1
Expert's answer
2022-01-17T10:47:02-0500

Balanced equation;

Fe2O3 + 2 Al → 2 Fe + Al2O3

Moles of Fe2O3 in 150g = "\dfrac{Mass}{Molar mass} =\dfrac{150g}{159.69g/mol}"

= 0.957 mol

Mole ratio of Al : Fe2O3 is 2:1


Moles of Al ="\dfrac{2}{1}x0.957 mol" = 1.915 mol


Mass of Al = Moles x RAM

= 1.915mol x 63.546g/mol

= 121.666g